# Learned Image Transmission with Hierarchical Variational Autoencoder

### v2

@@ III-D HJSCC with Feedback
III-E Masking and Length Information Reduction
A visualized example of our proposed masking strategy is shown in Fig. 5. Particularly, we employ the Swin Transformer blocks to implement our JSCC encoder, since we find that this architecture presents better robustness against channel noise. In this way, the shape of the output feature at the $l$-th layer, $\mathbf{r}_{l}$, can be denoted as $C\times H\times W$, which is the same as that of $\bm{\mu}_{l}$. We split the output of the Swin Transformer blocks, $\mathbf{r}_{l}$, into $HW$ sequences $\mathbf{r}_{l}^{o}$, for $o=1,2,\ldots,HW$, where the length of each $\mathbf{r}_{l}^{o}$ is $C$. Then, we include a rate-matching layer after the Swin Transformer blocks. It accepts two inputs, $\mathbf{r}_{l}$ and $\alpha p_{\mathbf{z}_{l}|\mathbf{z}_{<l}}$, and generates the masked channel symbol vector $\mathbf{s}_{l}$. Specifically, the length for each vector $\mathbf{s}_{l}^{o}$ will be constrained to $k_{l}^{o}=\sum_{c=1}^{C}-\alpha\log p_{z_{l}^{(o,c)}|\mathbf{z}_{<l}}(z_{l}^{(o,c)}|\mathbf{z}_{<l})$, which actually represents the summation of entropy along all the $C$ dimensions of $\mathbf{z}^{o}_{l}$. We adjust the length by generating a mask vector $\mathbf{m}_{l}^{o}$, which can be written as
@@ III-E Masking and Length Information Reduction
Figure 6: Detailed illustrations of the rate attention module.
@@
Figure 7: The end-to-end distortion performance versus the CBR over different datasets. The results are evaluated on (a) Kodak and (b) CLIC2022 datasets, at SNR $=10$ dB.
@@
Furthermore, as the JSCC encoder is required to encode the images into symbols of different numbers, we devise a rate attention module by incorporating the prior information in the encoding process. The detailed procedure of this module is depicted in Fig. 6. The rate attention involves two inputs, the latent representation $\bm{\mu}_{l}$ of shape $(H,W,C)$ and the prior information $p_{\mathbf{z}_{l}\mid\mathbf{z}_{<l}}$. We calculate the length value for different vectors in $\bm{\mu}_{l}$, obtaining in total $HW$ real length values. Then, we calculate the merged length values after the spatial grouping. As we allocate the same length value to the vectors within a patch, the merged value is calculated by averaging the real length values in a sample patch. The rate attention operation accepts two inputs. The first is the concatenated information from the representation and prior information, while the second is the concatenated matrix of the real length value vector and the merged length value vector. Through this operation, the index information can be fused to the encoding process, enabling the JSCC encoder block to adaptively adjust the encoding process.
IV Experiments
@@
For performance evaluation, we consider the pixel-wise peak signal-to-noise ratio (PSNR). In addition, the multi-scale structural similarity index (MS-SSIM) and the perceptual metric, learned perceptual image patch similarity (LPIPS) [34] are also included in the Appendix, which accounts for the nuances of human perception. We also use the BD-rate metric [35] to compute the average bit rate saving over all PSNRs. We quantify the performance by considering the following datasets of different resolutions with necessary preprocessing. Kodak [36]: The dataset consists of $24$ images of resolution $512\times 768$ or $768\times 512$. CLIC2022 Test [37]: The test set contains $30$ images up to size $1365\times 2048$.
@@
To verify the performance of our image transmission models, we compare them with a range of benchmarks. First, we consider emerging deep learning-based schemes, including DeepJSCC [9], SwinJSCC [30], and the nonlinear transform source-channel coding (NTSCC) [18]. Second, we compare our methods with separation-based schemes widely used in real transmission systems. These include powerful image codecs such as BPG, JPEG, and JPEG2000, combined with a practical LDPC code [38], labeled as “BPG + LDPC”, “JPEG + LDPC”, and “JPEG2000 + LDPC”, respectively. In addition to these hand-crafted image codecs, we also consider the learning-based image codecs combined with LDPC, [4] and [39]. For the feedback JSCC schemes, we consider the most advanced JSCCformer-f [13] and the classical DeepJSCC-f [12]. In our experiments, we test various schemes across different CBRs and SNRs under AWGN channels. We take the architecture in [25] as the backbone model, and the detailed architecture is presented in the Appendix.
Figure 8: The end-to-end distortion performance versus the SNR over different datasets. To ensure fairness, the average CBRs of all the methods are constrained to $0.0625$.
IV-A Comparisons and Results Analysis
@@
In Fig. 7, we evaluate the transmission performance at different CBRs under AWGN channels, with the test SNR set to $10$ dB. To ensure a reliable transmission link, we adopt $16$-order quadrature amplitude modulation (16QAM) combined with an LDPC rate of $2/3$, in accordance with the 3GPP standard. The results indicate that our proposed HJSCC significantly outperforms the fixed-rate schemes, DeepJSCC and SwinJSCC. Compared with NTSCC, the proposed scheme achieves comparable performance. The performance gap widens with increasing image resolution and CBR. Additionally, compared to hand-crafted schemes, the proposed method achieves substantially better PSNR performance. This demonstrates the potential of utilizing HJSCC in practical wireless communication systems.
Figure 9: PSNR performance of schemes with feedback link.
@@
Fig. 8 demonstrates the transmission performance across varying channel SNR levels. To ensure fairness, the CBR for these schemes is constrained to $0.0625$. For DeepJSCC and NTSCC, the training SNR equals the testing SNR to achieve optimal performance. For the separation-based methods, we test these schemes across different channel coding rates and modulation orders to determine the optimal settings [^1] Given a bpp value, the CBR $\rho$ can be calculated as $\rho=\frac{K}{C\times H\times W}=\frac{\text{bpp}}{C\times\log_{2}M\times R_{c}}=\frac{\text{bpp}}{C\times\log_{2}M\times R_{c}}$, where $M$ is the selected modulation order and $R_{c}$ denotes the channel coding rate. For example, $\rho=\text{bpp}/8$ when $16$QAM and rate $2/3$ are selected for SNR $=10$ dB., and then the image codec needs to compress the source with the resulted bits per pixel (bpp) value. Our results indicate that the proposed HJSCC significantly outperforms other schemes. Additionally, HJSCC achieves a substantial performance gain compared to the emerging method [39], with the performance gap becoming more pronounced at SNR levels above $10$ dB. Furthermore, separation-based systems are known to suffer from the cliff effect, where reliable transmission cannot be maintained when the channel coding and modulation schemes fail. In contrast, the proposed HJSCC provides a graceful degradation as the SNR decreases, demonstrating its robustness in varying channel conditions.
@@
Fig. 9 presents the PSNR performance of different schemes, where JSCCformer-f [13] and DeepJSCC-f [12] are JSCC schemes with feedback link. Compared with them, our proposed HJSCC shows much better PSNR performance, with a gain of about $1$ dB at high CBR region. Besides, the introduction of the feedback link also provides significantly larger gain, especially at high average CBR values. This stems from rate-adaptive capability of HJSCC-feedback as well as its generative formulation, making it the state-of-the-art JSCC schemes in the presence of a feedback link.
IV-B Ablation Studies
@@
In this work, we propose a spatial grouping strategy to reduce the transmission overhead to inform the receiver of the vector length for rate matching. To show the effectiveness, we report the performance on CBR saving. Particularly, we compare the CBRs for transmitting the channel symbols $\mathbf{s}$ and the vector length information $\hat{k}$ in Table I. All the models are optimized on ImageNet dataset. From the results, the overall CBR can be significantly reduced by the spatial merging strategy, at the cost of a slight performance degradation.
TABLE I: Ablation analysis on the spatial grouping strategy. The results are obtained on the Kodak dataset.

### v5

@@ Introduction
Masking and Length Information Reduction
A visualized example of our proposed masking strategy is shown in Fig. 5. Particularly, we employ the Swin Transformer blocks to implement our JSCC encoder, since we find that this architecture presents better robustness against channel noise. In this way, the shape of the output feature at the $l$-th layer, $\mathbf{r}_{l}$, can be denoted as $C\times H\times W$, which is the same as that of $\boldsymbol{\mu}_{l}$. We split the output of the Swin Transformer blocks, $\mathbf{r}_{l}$, into $HW$ sequences $\mathbf{r}_{l}^{o}$, for $o=1,2,\ldots,HW$, where the length of each $\mathbf{r}_{l}^{o}$ is $C$. Then, we include a rate-matching layer after the Swin Transformer blocks. It accepts two inputs, $\mathbf{r}_{l}$ and $\alpha p_{\mathbf{z}_{l}|\mathbf{z}_{<l}}$, and generates the masked channel symbol vector $\mathbf{s}_{l}$. Specifically, the length for each vector $\mathbf{s}_{l}^{o}$ will be constrained to $k_{l}^{o}=\sum_{c=1}^{C}-\alpha\log p_{z_{l}^{(o,c)}|\mathbf{z}_{<l}}(z_{l}^{(o,c)}|\mathbf{z}_{<l})$, which actually represents the summation of entropy along all the $C$ dimensions of $\mathbf{z}^{o}_{l}$. We adjust the length by generating a mask vector $\mathbf{m}_{l}^{o}$, which can be written as
@@
Figure 6: The end-to-end distortion performance versus over different datasets. The results are evaluated on (a) Kodak and (b) CLIC2022 datasets, at SNR $=10$ dB. (c) The end-to-end distortion performance on Kodak dataset versus the SNR.
@@
Furthermore, as the JSCC encoder is required to encode the images into symbols of different numbers, we devise a rate attention module by incorporating the prior information in the encoding process. The rate attention involves two inputs, the latent representation $\boldsymbol{\mu}_{l}$ of shape $(H,W,C)$ and the prior information $p_{\mathbf{z}_{l}\mid\mathbf{z}_{<l}}$. We calculate the length value for different vectors in $\boldsymbol{\mu}_{l}$, obtaining in total $HW$ real length values. Then, we calculate the merged length values after the spatial grouping. As we allocate the same length value to the vectors within a patch, the merged value is calculated by averaging the real length values in a sample patch. The rate attention operation accepts two inputs. The first is the concatenated information from the representation and prior information, while the second is the concatenated matrix of the real length value vector and the merged length value vector. Through this operation, the index information can be fused to the encoding process, enabling the JSCC encoder block to adaptively adjust the encoding process.
Experiments
@@
For performance evaluation, we consider the pixel-wise peak signal-to-noise ratio (PSNR) [^1] Codes are given in https://github.com/zhang-guangyi/HJSCC.. We quantify the performance by considering the following datasets of different resolutions with necessary preprocessing. Kodak (Kodak 1993): The dataset consists of $24$ images of resolution $512\times 768$ or $768\times 512$. CLIC2022 Test (Toderici et al. 2022): The test set contains $30$ images up to size $1365\times 2048$.
@@
To verify the performance of our image transmission models, we compare them with a range of benchmarks. First, we consider emerging deep learning-based schemes, including DeepJSCC (Bourtsoulatze et al. 2019), SwinJSCC (Yang et al. 2024), and the nonlinear transform source-channel coding (NTSCC) (Dai et al. 2022). Second, we compare our methods with separation-based schemes widely used in real transmission systems. These include powerful image codecs such as BPG, JPEG, and JPEG2000, combined with a practical LDPC code (Shahid and Yahampath 2013), labeled as “BPG + LDPC”, “JPEG + LDPC”, and “JPEG2000 + LDPC”, respectively. In addition to these hand-crafted image codecs, we also consider the learning-based image codecs combined with LDPC, (Ballé et al. 2018) and (Minnen, Ballé, and Toderici 2018). For the feedback JSCC schemes, we consider the most advanced JSCCformer-f (Wu et al. 2024a) and the classical DeepJSCC-f (Kurka and Gündüz 2020). In our experiments, we test various schemes across different CBRs and SNRs under AWGN channels. We take the architecture in (Duan et al. 2023) as the backbone model.
Comparisons and Results Analysis
@@
In Fig. 6 (a) and (b), we evaluate the transmission performance at different CBRs under AWGN channels, with the test SNR set to $10$ dB. To ensure a reliable transmission link, we adopt $16$-order quadrature amplitude modulation (16QAM) combined with an LDPC rate of $2/3$, in accordance with the 3GPP standard. The results indicate that our proposed HJSCC significantly outperforms the fixed-rate schemes, DeepJSCC and SwinJSCC. Compared with NTSCC, the proposed scheme achieves comparable performance. The performance gap widens with increasing image resolution and CBR. Additionally, compared to hand-crafted schemes, the proposed method achieves substantially better PSNR performance. This demonstrates the potential of utilizing HJSCC in practical wireless communication systems.
Figure 7: PSNR performance with feedback links.
@@
Fig. 6 (c) demonstrates the transmission performance across varying channel SNR levels. To ensure fairness, the CBR for these schemes is constrained to $0.0625$. For DeepJSCC and NTSCC, the training SNR equals the testing SNR to achieve optimal performance. For the separation-based methods, we test these schemes across different channel coding rates and modulation orders to determine the optimal settings [^2] Given a bpp value, the CBR $\rho$ can be calculated as $\rho=\frac{K}{C\times H\times W}=\frac{\text{bpp}}{C\times\log_{2}M\times R_{c}}=\frac{\text{bpp}}{C\times\log_{2}M\times R_{c}}$, where $M$ is the selected modulation order and $R_{c}$ denotes the channel coding rate. For example, $\rho=\text{bpp}/8$ when $16$QAM and rate $2/3$ are selected for SNR $=10$ dB., and then the image codec needs to compress the source with the resulted bits per pixel (bpp) value. Our results indicate that the proposed HJSCC significantly outperforms other schemes. Additionally, HJSCC achieves a substantial performance gain compared to the emerging method (Minnen, Ballé, and Toderici 2018), with the performance gap becoming more pronounced at SNR levels above $10$ dB. Furthermore, separation-based systems are known to suffer from the cliff effect, where reliable transmission cannot be maintained when the channel coding and modulation schemes fail. In contrast, the proposed HJSCC provides a graceful degradation as the SNR decreases, demonstrating its robustness in varying channel conditions.
@@
Fig. 7 presents the PSNR performance of different schemes, where JSCCformer-f (Wu et al. 2024a) and DeepJSCC-f (Kurka and Gündüz 2020) are JSCC schemes with a feedback link. Compared with them, our proposed HJSCC shows much better PSNR performance, with a gain of about $1$ dB at high CBR region. Besides, the introduction of the feedback link also provides significantly larger gain, especially at high average CBR values. This stems from rate-adaptive capability of HJSCC-feedback as well as its generative formulation, making it the state-of-the-art JSCC schemes in the presence of a feedback link. In addition, when the feedback is noisy, we can see a slight degradation in PSNR performance. Besides, the performance improves with the feedback link quality, where a higher feedback SNR can produce better PSNR performance.
Ablation Studies
@@
In this work, we propose a spatial grouping strategy to reduce the transmission overhead to inform the receiver of the vector length for rate matching. To show the effectiveness, we report the performance on CBR saving. Particularly, we compare the CBRs for transmitting the channel symbols $\mathbf{s}$ and the vector length information $\hat{k}$ in Table 1. All the models are optimized on ImageNet dataset. From the results, the overall CBR can be significantly reduced by the spatial merging strategy, at the cost of a slight performance degradation.
@@
Table 1: Ablation analysis on the spatial grouping strategy.
@@
Acknowledgments
The corresponding author is Yunlong Cai. This work was supported in part by the National Natural Science Foundation of China under Grant U22A2004, and in part by Zhejiang Provincial Key Laboratory of Information Processing, Communication and Networking (IPCAN), Hangzhou 310027, China.